OpenStudy (anonymous):

integral(sin^5(x)/(1+cos^2(x))dx super confused :(

OpenStudy (anonymous):

I don't even know how to begin starting this... \[\int\limits(\sin^5(x)/(1+\cos^2(x))dx\]

OpenStudy (anonymous):

im sorry, this ones' a little too difficult for me. sorry.

OpenStudy (anonymous):

I'm in Calculus II DX

OpenStudy (anonymous):

\[\int\limits(\frac{ (1-\cos^2(x))(1-\cos^2(x))(1-\cos^2(x) )}{ 1+\cos^2(x) }dx\] Am i doing this right....

OpenStudy (anonymous):

im sorry, im not sure, im taking calculus next semester. :(

OpenStudy (anonymous):

it needs to be done by substitution method...

OpenStudy (anonymous):

separate sin^5x into sin^4x * sinx and then suppose u=cosx do derivative then you get du=sinx dx after that it might be easy....all the trig function will go away...